OpenStudy (anonymous):

derivative of x^2 + y^2 - 8x - 6y + 15 = 0

OpenStudy (08surya):

2x+2ydy/dx-8-6dy/dx=0 dy/dx(2y-6)=-2x+8 dy/dx=(-2x+8)/(2y-6)